Ricky quits his job earning $150,000/year and converts a house that he owns into a museum for his large collection of antique clocks. Before this, Ricky had been renting the house out for $20,000/year. Ricky hired 2 employees for his museum at a total cost of $50,000/yr, and had to pay $10,000/yr for utilities.

His museum was a huge hit and brought in revenues of $125,000 the first year. One antique clock enthusiast even offered him $1,000,000 to purchase his entire collection, which he declined.

1. What is Ricky's accounting profit during the year? (accounting profit = total revenue - explicit costs)
2. What is Ricky's economic profit? Please explain. (economic profit = total revenue - explicit costs - implicit costs)
Answer #1

(1)

Accounting profit ($) = 125,000 - 50,000 - 10,000 = 65,000

(2)

Economic profit ($) = Revenue - explicit cost - Salary and rental income given up

= 125,000 - 50,000 - 10,000 - 150,000 - 20,000

= - 105,000

Note that offer of $100,000 given up is not a benefit given up for starting the business, so this is not a cost.